# Codes Correcting Transpositions of Consecutive Symbols

### v1

@@ 2.1 Construction over a general alphabet
Consider a codeword $\boldsymbol{x}\in\mathcal{C}_{q}(n)$. Let $\boldsymbol{y}$ be the string obtained after a pair of consecutive symbols in $\boldsymbol{x}$ have been swapped, meaning that $(y_{k},y_{k+1})=(x_{k+1},x_{k})$ for some $k\in\{1,\ldots,n-1\}$, and $y_{i}=x_{i}$ for all $i\notin\{k,k+1\}$. The index $k$ is the transposition location that needs to be determined in order to correct the error. The receiver computes
$c_{1}\coloneqq s_{1}$ $-\sum_{i=1}^{n}iy_{i}$ $=\sum_{i=1}^{n}ix_{i}$ $-\sum_{i=1}^{n}iy_{i}$ $=x_{k+1}-x_{k}\pmod{2q-1}$, (10a)
$c_{2}\coloneqq s_{2}$ $-\sum_{i=1}^{n}i^{2}y_{i}$ $=\sum_{i=1}^{n}i^{2}x_{i}$ $-\sum_{i=1}^{n}i^{2}y_{i}$ $=(2k+1)(x_{k+1}-x_{k})\pmod{p}$. (10b)
By computing $c_{1}$ in (10a), the receiver can uniquely determine the integer $x_{k+1}-x_{k}\in\{0,\pm 1,\ldots,\pm(q-1)\}$. If $c_{1}=0$, then $\boldsymbol{y}=\boldsymbol{x}$ and there is nothing to correct. Otherwise, the receiver can uniquely determine the error location $k$ from (10b), namely
$k=2^{-1}\big{(}c_{2}(x_{k+1}-x_{k})^{-1}-1\big{)}\pmod{p}$, (11)
and recover $\boldsymbol{x}$ by swapping the symbols $y_{k}$ and $y_{k+1}$. $\blacksquare$
Instead of a prime $p$ in (9), one could take any number $n^{\prime}\geqslant n$ that is mutually prime with all numbers in the range $2,\ldots,\max\{2,q-1\}$ (so that 22 and $x_{k+1}-x_{k}$ are guaranteed to have multiplicative inverses modulo $n^{\prime}$, see (11)). We have chosen to use a prime $p$ as it makes the definition a bit simpler while not affecting the cardinality of the resulting code in the asymptotic sense (see Theorem 2.3 ahead).
2.2 Construction over a binary alphabet
In this subsection we consider the case $q=2$. Although several constructions of single-transposition-correcting binary codes were reported earlier [22, 20, 16], we present here another construction that has the best known asymptotic scaling, equaling that obtained in [16], but that is somewhat more explicit than the one from [16].
Let $\mathcal{B}(m)\subseteq\mathbb{Z}_{2}^{m}$ be a binary code of length $m$ correcting one substitution error, i.e., having minimum Hamming distance at least 33. For a string $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$, let $\boldsymbol{x}_{\textnormal{even}}=(x_{2},x_{4},\ldots)$ be the string of its even-indexed symbols. Assume for simplicity that $n$ is even, and define
$\mathcal{C}^{\prime}(n)\coloneqq\left\{\boldsymbol{x}\in\mathbb{Z}_{2}^{n}\;\colon\boldsymbol{x}_{\textnormal{even}}\in\mathcal{B}(n/2),\;\;\sum_{i=1}^{n/2}i(x_{2i}-x_{2i-1})=s\pmod{2}\right\}$, (12)
where $s\in\{0,1\}$.
Theorem 2.2
Fix integers $n\geqslant 1$ and $s$ such that $\mathcal{C}^{\prime}(n)$ is nonempty. Then the code $\mathcal{C}^{\prime}(n)$ corrects one transposition.
Proof
Consider a codeword $\boldsymbol{x}\in\mathcal{C}^{\prime}(n)$, and let $\boldsymbol{y}$ be the string obtained after a transposition at location $k$ has been applied to $\boldsymbol{x}$ so that $(y_{k},y_{k+1})=(x_{k+1},x_{k})$ for some $k\in\{1,\ldots,n-1\}$, and $y_{i}=x_{i}$ for all $i\notin\{k,k+1\}$. The receiver, given $\mathcal{C}^{\prime}(n)$ and $\boldsymbol{y}$, needs to determine $k$. The string $\boldsymbol{y}_{\textnormal{even}}=({y}_{2},y_{4},\ldots,{y}_{n})$ is the result of applying (at most) one substitution error, i.e., bit-flip, to the string $\boldsymbol{x}_{\textnormal{even}}$. Recall that the latter is a codeword of a single-substitution-correcting code $\mathcal{B}(n/2)$. Hence, if $\boldsymbol{y}_{\textnormal{even}}\in\mathcal{B}(n/2)$, there was no substitution, meaning that the transposition had no effect on the transmitted string (because $x_{k}=x_{k+1}$), and there is nothing to correct. Otherwise, if $\boldsymbol{y}_{\textnormal{even}}\notin\mathcal{B}(n/2)$, the receiver will be able to find the erroneous symbol ${y}_{2j}$. Due to the nature of transposition errors, the transposition location $k$ is either $2j$ or 2j12j-1, and hence all that is left for the receiver to determine is whether $k$ is even or odd. To that end, the receiver computes
$c\coloneqq s-\sum_{i=1}^{n/2}i(y_{2i}-y_{2i-1})=\sum_{i=1}^{n/2}i(x_{2i}-x_{2i-1})-\sum_{i=1}^{n/2}i(y_{2i}-y_{2i-1})\pmod{2}$. (13)
It can be verified directly that, if k=2j1k=2j-1, then the sum in (13) equals $(k+1)(x_{k+1}-x_{k})$, and hence, since $x_{k+1}-x_{k}=\pm 1$, $c=0$. Similarly, if $k=2j$, the sum in (13) equals $(k+1)(x_{k}-x_{k+1})$, and hence $c=1$. Therefore, the calculated bit $c$ contains the information about the parity of $k$, and the transposition location can be computed as k=2j1+ck=2j-1+c. $\blacksquare$
2.3 Bounds
Theorem 2.3 below gives asymptotic bounds on the maximum possible cardinality of a code in $\mathbb{Z}_{q}^{n}$ correcting one transposition, namely $M^{\star}_{q}(n;1)$, and implies that the codes $\mathcal{C}_{q}(n)$ described in the previous subsection are asymptotically optimal in the sense that
$M^{\star}_{q}(n;1)=\Theta\!\left(|\mathcal{C}_{q}(n)|\right)=\Theta\Big{(}\frac{q^{n}}{n}\Big{)}.$ (14)
Before establishing the theorem, we state one auxiliary fact that will be needed in the proof.
Lemma 2
For fixed $\rho\in[0,1]$,
$\lim_{n\to\infty}\frac{1}{n}\log_{q}\!\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\;\colon\operatorname{run}(\boldsymbol{x})\leqslant\rho n\}\right|=\begin{cases}H_{q}(\rho),&0\leqslant\rho<1-q^{-1}\\ 1,&1-q^{-1}\leqslant\rho\leqslant 1\end{cases},$ (15)
where $H_{q}(\cdot)$ is the $q$-ary entropy function.
Proof
Let $S_{q}(n,r)$ denote the number of strings from $\mathbb{Z}_{q}^{n}$ having exactly $r$ runs of identical symbols, so that
$\left|\{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}\;\colon\operatorname{run}(\boldsymbol{x})\leqslant\rho n\}\right|=\sum_{r=1}^{\lfloor\rho n\rfloor}S_{q}(n,r).$ (16)
By a direct counting argument we find that
$S_{q}(n,r)=\binom{n-1}{r-1}q(q-1)^{r-1}$ (17)
and hence, by Stirling’s approximation,
$\lim_{n\to\infty}\frac{1}{n}\log_{q}S_{q}(n,\lfloor\rho n\rfloor)=H_{q}(\rho).$ (18)
$\max_{r\leqslant\rho n}S_{q}(n,r)\;\leqslant\;\sum_{r=1}^{\lfloor\rho n\rfloor}S_{q}(n,r)\;\leqslant\;\rho n\cdot\max_{r\leqslant\rho n}S_{q}(n,r),$ (19)
its exponential growth-rate (to base $q$) is
$\lim_{n\to\infty}\frac{1}{n}\log_{q}\sum_{r=1}^{\lfloor\rho n\rfloor}S_{q}(n,r)=\max_{0\leqslant\rho^{\prime}\leqslant\rho}H_{q}(\rho^{\prime}).$ (20)
That this is equal to the right-hand side of (15) follows from the fact that $H_{q}(\rho)$ is a concave function of $\rho\in[0,1]$ maximized at $\rho=1-q^{-1}$, the maximum being $H_{q}(1-q^{-1})=1$. $\blacksquare$
where $H(1-q^{-1}-\epsilon)<1$. A very useful consequence of this fact is that one can disregard such strings in the asymptotic analysis of optimal codes. Namely, by noticing that any code correcting 22 substitutions, i.e., having minimum Hamming distance >4>\!4, is also capable of correcting one transposition (because a transposition affects two symbols in the original string), and by using the best known such codes, it follows that $M^{\star}_{q}(n;1)=\Omega(\frac{q^{n}}{n^{4}})$ and hence
$M^{\star}_{q}(n;1)=\Theta\!\left(q^{n+o(n)}\right).$ (22)
Relations (21) and (22) imply that, for an optimal single-transposition-correcting code, throwing out its codewords that have at most $(1-q^{-1}-\epsilon)n$ runs affects its cardinality only by an asymptotically negligible amount.
Theorem 2.3
For any fixed $q\geqslant 2$, as $n\to\infty$,
$\frac{q^{n}}{(2q-1)\,n}\;\lesssim\;M^{\star}_{q}(n;1)\;\lesssim\;\frac{q^{n+1}}{(q-1)\,n}.$ (23)
In the binary case ($q=2$), the lower bound can be improved to $\frac{2^{n}}{n}$.
Proof
The lower bound in (23) follows from the construction presented in Section 2.1. Namely, by the pigeon-hole principle there exist parameters $s_{1}\in\{1,\ldots,2q-1\}$ and $s_{2}\in\{1,\ldots,p\}$ such that the code $\mathcal{C}_{q}(n)$ has at least $\frac{q^{n}}{(2q-1)\,p}$ codewords (see (9)). Since the smallest prime $p$ greater than or equal to $n$ satisfies $p=n+o(n)$ when $n\to\infty$ [4, Section 1.4.1], the lower bound in (23) follows.
Similarly, the lower bound of $\frac{2^{n}}{n}$ for the binary case follows from the construction presented in Section $2.$ (We note that this bound was also reported in [16, Theorem 11].) Namely, by the pigeon-hole principle, for at least one value of $s\in\{0,1\}$ the code $\mathcal{C}^{\prime}(n)$ has at least $\frac{|\mathcal{B}(n/2)|\cdot 2^{n/2}}{2}$ codewords. The result then follows by recalling that there exist binary single-substitution-correcting codes $\mathcal{B}(m)$ (e.g., Hamming codes) satisfying $|\mathcal{B}(m)|\gtrsim\frac{2^{m}}{m}$.
The upper bound in (23) is obtained by a packing argument. Let $\mathcal{C}^{\star}_{q}(n)$ be an optimal code correcting one transposition, meaning that $|\mathcal{C}^{\star}_{q}(n)|=M^{\star}_{q}(n;1)$. Recall that $B(\boldsymbol{x};1)$ denotes the set of all strings that can be obtained by applying at most one transposition to a string $\boldsymbol{x}$, and note that
$|B(\boldsymbol{x};1)|=\operatorname{run}(\boldsymbol{x}).$ (24)
This is because only the transpositions occurring at the boundaries of consecutive runs in $\boldsymbol{x}$ will actually alter $\boldsymbol{x}$, and hence there are $\operatorname{run}(\boldsymbol{x})-1$ strings different from $\boldsymbol{x}$ that can be obtained in this way. Accounting for $\boldsymbol{x}$ itself, (24) follows. We further have
$\sum_{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n)\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n}|B(\boldsymbol{x};1)|\;\leqslant\;\sum_{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n)}|B(\boldsymbol{x};1)|\;\leqslant\;|\mathbb{Z}_{q}^{n}|\;=\;q^{n},$ (25)
where the second inequality follows from the fact that the sets $B(\boldsymbol{x};1)$ and $B(\boldsymbol{x}^{\prime};1)$ are disjoint for any two distinct codewords $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathcal{C}^{\star}_{q}(n)$ (see (2)). From (24) and (25) we then obtain
$(1-q^{-1}-\epsilon)n\cdot\left|\left\{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n)\;\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n\right\}\right|\;<\;q^{n}.$ (26)
From the discussion after the proof of Lemma 2 we know that we can safely ignore the codewords with at most $(1-q^{-1}-\epsilon)n$ runs in the asymptotic analysis, meaning that
$\left|\left\{\boldsymbol{x}\in\mathcal{C}^{\star}_{q}(n)\;\colon\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n\right\}\right|\;\sim\;\left|\mathcal{C}^{\star}_{q}(n)\right|\;=\;M^{\star}_{q}(n;1).$ (27)
Now (26) and (27) imply
$\left(1-q^{-1}-\epsilon\right)n\cdot M^{\star}_{q}(n;1)\;\lesssim\;q^{n}.$ (28)
Since (28) holds for any $\epsilon>0$, the upper bound from (23) is established. $\blacksquare$
Consequently, as $n\to\infty$,
$M^{\star}_{q}(n;1)\;=\;\Theta\Big{(}\frac{q^{n}}{n}\Big{)},$ (29)
or, written differently,
$\log_{q}M^{\star}_{q}(n;1)\;=\;n-\log_{q}n+{\mathcal{O}}(1).$ (30)
Resilience to one transposition error requires $\log_{q}n+{\mathcal{O}}(1)$ redundant symbols.
3 Codes correcting $t=\textnormal{const}$ transpositions
In this section we present bounds on the cardinality of optimal codes in $\mathbb{Z}_{q}^{n}$ correcting $t$ transpositions (Theorem 3.1 ahead). The upper bound generalizes to arbitrary $t$ the bound stated in Theorem 2.3 for the special case $t=1$.
Recall that $B(\boldsymbol{x};t)$ denotes the set of all strings that can be obtained by applying up to $t$ transpositions to the string $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$, and $\bar{B}(\boldsymbol{x};t)$ denotes a ball (with respect to distance $d(\cdot,\cdot)$) of radius $t$ around $\boldsymbol{x}$. We first state bounds on these quantities that will be useful in what follows.
Lemma 3
For every $q\geqslant 2$, $n\geqslant 1$, $t\geqslant 1$, and $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$,
$\sum_{u=0}^{t}\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor}{u}\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor-2u-1}{t-u}\;\leqslant\;|B(\boldsymbol{x};t)|\;\leqslant\;\sum_{s=0}^{t}\binom{n-s}{s}$. (31)
Proof
To derive a lower bound on the cardinality of $B(\boldsymbol{x};t)$, consider the odd-numbered runs (first, third, etc.) in $\boldsymbol{x}$. Every transposition in $\boldsymbol{x}$ occurs either at the left or at the right boundary of an odd-numbered run, and so every pattern of exactly $t$ transpositions can be described by assuming that $u$ transpositions occur at the right boundary of such a run in the first phase (i.e., we select $u$ odd-numbered runs and swap their right-most symbols with the left-most symbols of the even-numbered runs immediately to their right), after which tut-u transpositions occur at the left boundary of such a run in the second phase, for some $u=0,1,\ldots,t$. The former can happen in $\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor}{u}$ ways, and the latter in at least $\binom{\lfloor\operatorname{run}(\boldsymbol{x})/2\rfloor-2u-1}{t-u}$ ways, implying the left-hand inequality in (31). (In the second phase we exclude the first run, because it has no boundary on the left. Also, if the $i$’th run was selected for the transposition in the first phase, we exclude in the second phase both the $i$’th run, because it is possible that it is of length 11 whence its left and right boundaries coincide, and the $(i+2)$’th run, because it is possible that the $(i+1)$’th run in between them was of length 11. Hence the 2u1-2u-1 term.)
To prove the stated upper bound on the cardinality of $B(\boldsymbol{x};t)$, it suffices to show that there are at most $\binom{n-s}{s}$ strings that can be obtained by applying exactly $s$ transpositions to $\boldsymbol{x}$. Since the transposition locations $k_{1},\ldots,k_{s}$ belong to $\{1,\ldots,n-1\}$ and satisfy $k_{j+1}>k_{j}+1$, every set of $s$ transposition locations corresponds uniquely to a binary string with $s$ 11’s and nsn-s 0’s such that there is at least one 0 after every 11. The number of such strings can be found by counting the number of ways to distribute n2sn-2s 0’s in $s+1$ blocks (where the blocks are separated by pairs of symbols $1,0$), which is $\binom{(n-2s)+(s+1)-1}{(s+1)-1}=\binom{n-s}{s}$. $\blacksquare$
Lemma 4
For every $q\geqslant 2$, $n\geqslant 1$, $t\geqslant 1$, and $\boldsymbol{x}\in\mathbb{Z}_{q}^{n}$,
$|\bar{B}(\boldsymbol{x};t)|$ $\;\leqslant\;\sum_{v=0}^{2t/3}\binom{2(t-v)}{v}|B(\boldsymbol{x};t-v)|$ (32a)
$\;\leqslant\;\Big{(}\frac{2t}{3}+1\Big{)}\max_{v}\binom{2(t-v)}{v}|B(\boldsymbol{x};t-v)|.$ (32b)
Proof
We can think of $\bar{B}(\boldsymbol{x};t)$ as the set of all strings that can be obtained through a two-step process: first apply up to $u$ transpositions to $\boldsymbol{x}$, resulting in a string $\boldsymbol{y}$, and then apply $v$ transpositions to $\boldsymbol{y}$, where $u+v=t$.
Taking $u=t$ in the first step (and hence $v=0$ in the second), we get exactly $B(\boldsymbol{x};t)$, i.e., the set of strings that can be obtained from $\boldsymbol{x}$ with at most $t$ transpositions. This is the summand $v=0$ in (32a).
Consider now the case u=t1u=t-1, and let $\boldsymbol{y}\in B(\boldsymbol{x};t-1)$ be a string obtained from $\boldsymbol{x}$ in the first step. In the second step we are allowed to choose $v=1$ pair of consecutive symbols in $\boldsymbol{y}$ to swap. If this pair is disjoint from all the pairs chosen in the first step, then the resulting string $\boldsymbol{z}$ can equivalently be obtained directly from $\boldsymbol{x}$ via up to $t$ disjoint transpositions. In other words, $\boldsymbol{z}\in B(\boldsymbol{x};t)$ and this string was therefore already accounted for in the first case described in the previous paragraph. Also, if the pair chosen in the second step is identical to a pair chosen in the first step, then the transposition from the second step will annul the transposition from the first step and, again, $\boldsymbol{z}\in B(\boldsymbol{x};t)$, i.e., we are back to the previous case. Therefore, in the second step we can, without loss of generality, consider only pairs of symbols that partially overlap with a pair chosen in the first step, either on its left or on its right symbol. (E.g., suppose that $(a,b,c,d)$ is a substring of $\boldsymbol{x}$, and that the pair $(b,c)$ was chosen in the first step, so that the corresponding substring of $\boldsymbol{y}$ is $(a,c,b,d)$. Then the pairs in $\boldsymbol{y}$ that partially overlap with the pair chosen in the first step, one of which we can choose in the second step, are $(a,c)$ and $(b,d)$.) Since at most $u$ pairs were chosen in the first step, there are at most $2u$ pairs to select one from in the second step. In conclusion, there are $|B(\boldsymbol{x};u)|$ different strings $\boldsymbol{y}$ that can be obtained in the first step, and for each such $\boldsymbol{y}$ there are at most $\binom{2u}{1}=2u$ different strings $\boldsymbol{z}$ that can be obtained in the second step.
The analysis in the general case is the same: there are $|B(\boldsymbol{x};u)|$ different strings $\boldsymbol{y}$ that can be obtained in the first step, and for each such $\boldsymbol{y}$ there are at most $\binom{2u}{v}$ different strings $\boldsymbol{z}$ that can be obtained in the second step. $\blacksquare$
Theorem 3.1
For any fixed $q\geqslant 2$ and $t\geqslant 1$, as $n\to\infty$,
$\frac{(2t)!\,q^{n}}{(4t/3+1)\,n^{2t}}\;\lesssim\;M^{\star}_{q}(n;t)\;\lesssim\;\frac{t!\,q^{n+t}}{(q-1)^{t}\,n^{t}}.$ (33)
In the binary case ($q=2$), the lower bound can be improved to $c_{t}\frac{2^{n+t}}{n^{t}}$, where $c_{1}=\frac{1}{2}$, $c_{2}=\frac{1}{3}$, and $c_{t}=\frac{1}{2t+1}$ for $t\geqslant 3$.
Proof
As noted in Section 1.1, the distance function $d(\cdot,\cdot)$ is not a metric, and hence $\bar{B}(\boldsymbol{x};t)$ is not a metric ball. Nonetheless, the same argument used to obtain the Gilbert–Varshamov bound in classical settings [8, 28] can be applied here as well. Namely, by a greedy procedure one is guaranteed to find a collection of strings $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ satisfying the following two conditions:
@@ 3 Codes correcting $t=\textnormal{const}$ transpositions
The first condition implies, by Proposition 1, that $\mathcal{C}$ is a code correcting $t$ transpositions. The second condition, stating that balls of radius $2t$ around the codewords cover the whole space, implies that $\sum_{\boldsymbol{x}\in\mathcal{C}}|\bar{B}(\boldsymbol{x};2t)|\geqslant q^{n}$, and hence
$|\mathcal{C}|\;\geqslant\;\frac{q^{n}}{\max_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};2t)|}.$ (34)
By Lemmas 3 and 4, we have
$\max_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};2t)|\;\leqslant\;\Big{(}\frac{4t}{3}+1\Big{)}\max_{v}\binom{2(2t-v)}{v}\sum_{s=0}^{2t-v}\binom{n-s}{s}$. (35)
Since $\binom{n-s}{s}\sim\frac{n^{s}}{s!}$ when $n\to\infty$, the dominant term on the right-hand side of (35) is obtained for $v=0$ and $s=2t$, implying
$\max_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};2t)|\;\lesssim\;\Big{(}\frac{4t}{3}+1\Big{)}\frac{n^{2t}}{(2t)!}.$ (36)
By (34) and (36), we have established the existence of a family of $t$-transposition-correcting codes of cardinality $\gtrsim\frac{(2t)!\,q^{n}}{(4t/3+1)\,n^{2t}}$.
The lower bound of $c_{t}\frac{2^{n+t}}{n^{t}}$ for the binary case was derived in [16, Theorem 11] for the stronger model with successive transpositions, and is therefore automatically valid for the model with disjoint transpositions (see Section 1.2).
In order to show the upper bound in (33), an asymptotic lower bound on the cardinality of $B(\boldsymbol{x};t)$ will be needed. Using Lemma 3, for strings satisfying $\operatorname{run}(\boldsymbol{x})>(1-q^{-1}-\epsilon)n$, $|B(\boldsymbol{x};t)|$ can be asymptotically lower-bounded by
$\gtrsim\;\sum_{u=0}^{t}\binom{(1-q^{-1}-\epsilon)n/2}{u}\binom{(1-q^{-1}-\epsilon)n/2}{t-u}$ (37a)
$\sim\;\sum_{u=0}^{t}\frac{(1-q^{-1}-\epsilon)^{u}\,n^{u}}{2^{u}\,u!}\frac{(1-q^{-1}-\epsilon)^{t-u}\,n^{t-u}}{2^{t-u}\,(t-u)!}$ (37b)
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{2^{t}}\,\sum_{u=0}^{t}\frac{1}{u!\;(t-u)!}$ (37c)
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{2^{t}\,t!}\,\sum_{u=0}^{t}\binom{t}{u}$ (37d)
$=\;\frac{(1-q^{-1}-\epsilon)^{t}\,n^{t}}{t!}.$ (37e)
(In (37b) we used the fact that $\binom{n}{s}\sim\frac{n^{s}}{s!}$ when $s$ is fixed and $n\to\infty$.) The rest of the proof is similar to that of Theorem 2.3. One first notes that, by (21), disregarding all codewords having at most $(1-q^{-1}-\epsilon)n$ runs incurs no penalty in terms of the asymptotic scaling. Further, by (2), the sets $B(\boldsymbol{x};t)$ centered at codewords of an arbitrary code $\mathcal{C}$ correcting $t$ transpositions do not overlap, and hence $\sum_{\boldsymbol{x}\in\mathcal{C}}|B(\boldsymbol{x};t)|\leqslant|\mathbb{Z}_{q}^{n}|$. These facts, together with (37), imply
$M^{\star}_{q}(n;t)\cdot\frac{1}{t!}\left(1-q^{-1}-\epsilon\right)^{\!t}n^{t}\;\lesssim\;q^{n}.$ (38)
Since this relation holds for any $\epsilon>0$, the upper bound in (33) follows. $\blacksquare$
Therefore, attaining resilience to $t$ transpositions requires at least $t\log_{q}n+{\mathcal{O}}(1)$ and at most $2t\log_{q}n+{\mathcal{O}}(1)$ redundant symbols. Based on what is known in the special case $t=1$ (see (29)), as well as in the case of general $t$ and $q=2$, namely $M^{\star}_{2}(n;t)=\Theta(\frac{2^{n}}{n^{t}})$, it is reasonable to conjecture that
i.e., that the redundancy of $t\log_{q}n+{\mathcal{O}}(1)$ is achievable for any fixed alphabet size $q$. At present, however, we do not have a lower bound on the cardinality of optimal codes that would imply this result when $q>2$ and $t>1$.

### v2

@@ 3.1 Generalized Gilbert–Varshamov bound
As noted in Section 1.1, the distance function $d(\cdot,\cdot)$ is not a metric, and hence $\bar{B}(\boldsymbol{x};r)$ is not a metric ball. Nonetheless, the same argument used to obtain the Gilbert–Varshamov bound in classical settings [11, 31, 13] can be applied in our setting as well. Namely, for any given $t$, by using a greedy procedure one is guaranteed to find a collection of strings $\mathcal{C}\subseteq\mathbb{Z}_{q}^{n}$ satisfying the following two conditions:
@@
The first condition implies, by Proposition 1, that $\mathcal{C}$ is a code correcting $t$ transpositions. The second condition, stating that balls of radius $2t$ around the codewords cover the whole space, implies that $\sum_{\boldsymbol{x}\in\mathcal{C}}|\bar{B}(\boldsymbol{x};2t)|\geqslant q^{n}$. Therefore, it follows from the argument in [13] that
${M}_{q}^{\star}(n;t)\;\geqslant\;\frac{q^{n}}{\frac{1}{q^{n}}\sum_{\boldsymbol{x}\in\mathbb{Z}_{q}^{n}}|\bar{B}(\boldsymbol{x};2t)|}\;=\;\frac{q^{2n}}{\bar{T}_{q}(n;2t)},$ (31)
where $\bar{T}_{q}(n;r)$ is the “total ball size” of radius $r$ (see (16)).
If we denote the exponential growth rate (to base 22) of $\bar{T}_{q}(n;\lfloor\rho n\rfloor)$ by
$\bar{\beta}_{q}(\rho)\coloneqq\lim_{n\to\infty}\frac{1}{n}\log_{2}\bar{T}_{q}(n;\lfloor\rho n\rfloor),$ (32)
then (31) implies that
${R}^{\star}_{q}(\tau)\;\geqslant\;2\log_{2}(q)-\bar{\beta}_{q}(2\tau).$ (33)
From Proposition 6 and the fact that $\lim_{n\to\infty}\frac{1}{n}\log_{2}\!\binom{\alpha n}{\gamma n}=\alpha H(\frac{\gamma}{\alpha})$ one can obtain the exponential growth rate of $T^{\textnormal{e}}_{q}(n;\lfloor\rho n\rfloor)$, namely
$\lim_{n\to\infty}\frac{1}{n}\log_{2}T^{\textnormal{e}}_{q}(n;\lfloor\rho n\rfloor)=\beta^{\textnormal{e}}_{q}(\rho),$ (34)
where $\beta^{\textnormal{e}}_{q}(\rho)$ is given by (27). Note that, since the maximum number of disjoint transpositions is $\lfloor n/2\rfloor$, the domain of this function is $0,\frac{1}{2}]$. Then from (17) one obtains the exponential growth rate of $T_{q}(n;\lfloor\rho n\rfloor)$, namely
$\lim_{n\to\infty}\frac{1}{n}\log_{2}T_{q}(n;\lfloor\rho n\rfloor)=\beta_{q}(\rho),$ (35)
where $\beta_{q}(\rho)$ is given by (29). Finally, Proposition 5 implies that
$\bar{T}_{q}(n;r)\;\leqslant\;\Big(\frac{2r}{3}+1\Big)\max_{\lceil\frac{r}{3}\rceil\leqslant u\leqslant\min\{r,\lfloor\frac{n}{2}\rfloor\}\}}\binom{2u}{r-u}T_{q}(n;u),$ (36)
which further implies that, for any $q\geqslant 2$ and $\rho\in[0,1]$,
$\bar{\beta}_{q}(\rho)\;\leqslant\;\max_{\frac{\rho}{3}\leqslant\lambda\leqslant\min\{\rho,\frac{1}{2}\}}\left[2\lambda H\Big(\frac{\rho-\lambda}{2\lambda}\Big)+\beta_{q}(\lambda)\right].$ (37)
The theorem follows from (33) and (37). $\blacksquare$
The bound from Theorem 3.1 is plotted in Figure 1 for an alphabet of size $q=4$. For comparison, we also plot the bound
${R}^{\star}_{q}(\tau)\;\geqslant\;\begin{cases}\log_{2}(q)-2\tau\log_{2}(q-1)-H(2\tau),&0\leqslant\tau\leqslant\frac{q-1}{2q}\\ 0,&\tau>\frac{q-1}{2q}\end{cases}.$ (38)
The expression on the right-hand side of (38) is the familiar Gilbert–Varshamov bound for codes correcting $\tau n$ substitutions, i.e., codes of minimum Hamming distance $2\tau n$ [15, Section 2.10.6]. Recall from Section 1.3 that any code correcting $t$ substitutions can be used to construct a code of the same cardinality correcting $t$ transpositions.